OpenStudy (anonymous):

Use the matrix tool to find the solution to this system of equations. 2x + 4y + 2z = 16 3x + 2y + 2z = 13 -4x + 5y + 3z = 15 z = What is the z-value of the solution? A. 4 B. 3 C. 2 D. 5

OpenStudy (anonymous):

What's the "matrix tool"?

Parth (parthkohli):

@estudier Solve the equation by any method. You just have to get the answer.

OpenStudy (anonymous):

Soo what is it

Parth (parthkohli):

Okay. -x + 2y = 3

Parth (parthkohli):

-4x + 8y = 12

Parth (parthkohli):

13y + 3z = 27

Parth (parthkohli):

z = 3